OpenStudy (anonymous):

Which table best represents the steps to solve the equation x over 2 + 1 = -5?

OpenStudy (anonymous):

Step Justification Step 1: x over 2 = -5 + 1 The same number +1 can be added to both sides of the equation Step 2: x over 2 = -4 Adding 1 and -5 on the right hand side Step 3: x = -4 • 2 The same number 2 can be multiplied on both sides of the equation Step 4: x = -8 Multiplying -4 and 2 on the right hand side Step Justification Step 1: x over 2 = -1 - 5 The same number - 1 can be added to both sides of the equation Step 2: x over 2 = -6 Adding -1 and -5 on the right hand side Step 3: x = -6 - 2 The same number 2 can be subtracted on both sides of the equation Step 4: x = -8 Adding -6 and -2 Step Justification Step 1: x over 2 = -1 - 5 The same number -1 can be added to both sides of the equation Step 2: x over 2 = -6 Adding -1 and -5 on the right hand side Step 3: x = -6 • 2 The same number 2 can be multiplied on both sides of the equation Step 4: x = -12 Multiplying -6 and 2 on the right hand side Step Justification Step 1: x over 2 = -1 - 5 The same number -1 can be added to both sides of the equation Step 2: x over 2 = -6 Adding -1 and -5 on the right hand side Step 3: x = -6 ÷ 2 The same number 2 can be divided on both sides of the equation Step 4: x = -3 Dividing -6 by 2 on the right hand side